OpenStudy (cece4918):

x varies inversely with y and x = 6 when y = 10. What is the value of y when x = 8? y=4/15 y=4.8 y=7.5 this is the correct answer but i don't understand why can someone please help me understand why y=40/3

OpenStudy (learner):

\(x\) varies inversely with \(y\) means that: \(\color{blue}{\displaystyle y=\frac{k}{x}}\)

OpenStudy (learner):

You are given that when \(x=6\), then \(y=10\). So, plug that in ...

OpenStudy (cece4918):

Thank you so much this helps a lot!

OpenStudy (learner):

\(\color{red}{\displaystyle x=6}\) \(\color{green}{\displaystyle y=10}\) \(\color{blue}{\displaystyle \color{green}{y}=\frac{k}{\color{red}{x}}\quad \Longrightarrow \quad \quad \color{green}{10}=\frac{k}{\color{red}{6}}}\)

OpenStudy (learner):

You can solve for \(k\) (by multiplying times \(6\) on both sides).

OpenStudy (learner):

\(\color{blue}{\displaystyle 10\color{red}{\times 6}=\frac{k}{6}\color{red}{\times 6}}\) \(\color{blue}{\displaystyle 60=k}\)

OpenStudy (learner):

So, we found the \(k\).

OpenStudy (learner):

Therefore, you have the following inverse variation equation: \(\color{blue}{\displaystyle y=\frac{60}{x}}\)

OpenStudy (learner):

Now, (in the question) you are asked to find the value of \(y\), when \(x=8\). To do this, plug in \(8\) instead of \(x\), and solve for \(y\).

OpenStudy (cece4918):

thank you!

OpenStudy (learner):

yw